### v1

# Eccentric and cool: a high-spectral resolution view of

11institutetext: Aix Marseille Univ, CNRS, CNES, LAM, Marseille, France
11email: allan.denis@lam.fr
22institutetext: Max-Planck-Institut f¨ur Astronomie, K¨onigstuhl 17, 69117 Heidelberg, Germany 33institutetext: Univ. Grenoble Alpes, CNRS, IPAG, F-38000 Grenoble, France 44institutetext: Laboratoire J.L. Lagrange, Université Côte d’Azur, Observatoire de la côte d’Azur, CNRS, 06304 Nice, France 55institutetext: LIRA, Observatoire de Paris, Université PSL, Sorbonne Université, Université de Paris, 5 place Jules Janssen, 92195 Meudon, France 66institutetext: Department of Physics & Astronomy, John Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA 77institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 88institutetext: Leiden Observatory, Leiden University, Einsteinweg 55, 2333 CC Leiden, The Netherlands 99institutetext: NASA-Goddard Space Flight Center, Greenbelt, MD 20771, USA 1010institutetext: Department of Astronomy, Stockholm University, AlbaNova University Center, 10691 Stockholm, Sweden 1111institutetext: Institute for Astrophysics und Geophysik, Georg-August University, Friedrich-Hund-Platz 1, 37077 Göttingen, Germany 1212institutetext: Universitäts-Sternwarte, Ludwig-Maximilians-Universität München, Scheinerstraße 1, 81679 München, Germany 1313institutetext: Exzellenzcluster Origins, Boltzmannstraße 2, 85748 Garching, Germany 1414institutetext: Instituto de Estudios Astrofísicos, Facultad de Ingeniería y Ciencias, Universidad Diego Portales, Av. Ejército 441, Santiago, Chile 1515institutetext: European Southern Observatory, Alonso de Cordova 3107, Vitacura, Santiago, Chile 1616institutetext: Millennium Nucleus on Young Exoplanets and their Moons 1717institutetext: Department of Physics and Astronomy, University of Texas-San Antonio, San Antonio, TX, USA 1818institutetext: Center for Advanced Instrumentation, Durham University, Durham, DH1 3LE, United Kindgom 1919institutetext: Dept. of Astrophysics, University of Oxford, Keble Road, Oxford, OX1 3RH, UK 2020institutetext: Academia Sinica, Institute of Astronomy and Astrophysics, 11F Astronomy-Mathematics Building, NTU/AS campus, No. 1, Section 4, Roosevelt Rd., Taipei 10617, Taiwan 2121institutetext: European Southern Observatory (ESO), Karl-Schwarzschild-Str. 2, 85748 Garching, Germany 2222institutetext: Physics & Astronomy Dpt, University of Exeter, Exeter, EX4 4QL, UK 2323institutetext: Institute for Astronomy, University of Hawaii at Manoa, Honolulu, HI 96822, USA
Eccentric and cool: a high-spectral resolution view of
51 Eri b with VLT/HiRISE [^1] Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programmes ID 112.25FU, 114.2712, and 115.284P
A. Denis    A. Vigan    G. Chauvin    A. Lacquement    H. Beust    M. Ravet    J. Costes    A. Radcliffe    S. Martos    W. Balmer    T. Stolker    P. Palma-Bifani    B. Rajpoot    A. Simonnin    S. Petrus    T. Forveille    M. Janson    A. Reiners    N. Godoy    D. Cont    L. Nortmann    K. Hoy    A. Zurlo    H. Anwand-Heerwart    Y. Charles    A. Costille    M. El Morsy    J. Garcia    M. Houllé    M. Lopez    G. Murray    E. Muslimov    G. P. P. L. Otten    J. Paufique    M. Phillips    U. Seemann    A. Viret    G. Zins
(Received 05 January 2026; accepted 31 January 2026)
Discovered almost 10 years ago, the giant planet 51 Eridani b is one of the least separated ($\approx$0.2′′) and faintest (J $\approx$ 19.74 mag) directly imaged exoplanets known to date. Its atmospheric properties have been thoroughly investigated through low- and medium-resolution spectroscopic observations, enabling robust characterization of the planet’s bulk parameters. However, the planet’s intrinsically high contrast renders high-resolution spectroscopic observations difficult, despite their potential to yield key measurements essential for a more comprehensive characterization. This study seeks to constrain the planet’s radial velocity, enabling a full 3D orbital solution when integrated with previous measurements. We have obtained 4 high-contrast high-resolution (R $\approx$ 140 000) spectroscopic datasets of the planet, collected over a two-year interval with the HiRISE visitor instrument at the VLT to derive the planet’s radial velocity. Using self-consistent models of atmosphere, we were able to derive the radial velocity of the planet at each of the 4 epochs. These radial velocity measurements were then used in combination with all existing relative astrometry in order to constrain the orbit of the planet.Our radial velocity measurements allow us to break the degeneracy along the line of sight, making it now possible the unambiguous interpretation of the phase curve of the companion. We further constrain the orbital parameters, particularly the eccentricity, for which we derive e = $0.55_{-0.07}^{+0.03}$. The relatively high eccentricity indicates that the system has experienced dynamical interactions induced by an external perturber. We place constraints on the mass and semi-major axis of a hypothetical, unseen outer planet capable of producing the observed high eccentricities.
Key Words.:
Instrumentation: high angular resolution – Instrumentation: adaptive optics – Instrumentation: spectrographs – Techniques: high-angular resolution – Techniques: spectroscopy – Infrared: planetary systems
1 Introduction
Since the first spatially resolved exoplanet detection (Chauvin et al. 2004), direct imaging has enabled the discovery and characterization of companions at wide separations, probing the outer regions of exoplanetary systems (Bowler 2016). Population-level studies have provided initial constraints on their demographics and formation pathways (Nielsen et al. 2019; Vigan et al. 2021), and, when combined with other detection techniques, direct imaging contributes to a global view of planetary system architectures across a broad range of masses and separations (Bonfils et al. 2013; Batalha 2014; Bowler et al. 2020).
Eccentricity, for instance, is an efficient indicator of planetary formation and dynamical evolution. While planets forming in protoplanetary disks are expected to form on circular orbits, they can develop non-zero eccentricities through dynamical processes such as planet–planet scattering (Ford and Rasio 2008), Kozai–Lidov oscillations (Kozai 1962; Lidov 1962), planet–disk interactions (Li and Lai 2023), or stellar flybys (Pfalzner et al. 2024). A population-level dichotomy exists between brown dwarf and giant planet companions, with brown dwarfs peaking at high eccentricities ($e\approx 0.6$$0.9$) and giant planets typically exhibiting low eccentricities (Bowler et al. 2020). The giant planet 51 Eridani b (hereafter 51 Eri b) is an exception, as recent studies suggest a potentially high eccentricity (Rosa et al. 2020; Dupuy et al. 2021; Balmer et al. 2025), indicative of past dynamical interactions.
Discovered in 2014 by the Gemini Planet Imager Exoplanet Survey (Macintosh et al. 2015), 51 Eri b is orbiting an F0 star that is a member of the $\beta$ Pictoris moving group ($24\pm 3$ Myr; Bell et al. 2015). Updated stellar parameters yield $R_{\star}=1.45\pm 0.02,R_{\odot}$ and $M_{\star}=1.550\pm 0.006,M_{\odot}$, with an age of $23\pm 1$ Myr (Elliott et al. 2024). The system also hosts the wide, co-moving M-dwarf binary GJ 3305 AB at a projected separation of $\sim$2000,au (Feigelson et al. 2006; Kasper et al. 2007).
[…]
In this work, we present nearly two years of high-resolution ($R=140,000$) direct spectroscopic observations of 51 Eri b obtained with the HiRISE instrument (Vigan et al. 2024), combining SPHERE (Beuzit et al. 2019) and CRIRES (Dorn et al. 2023). We describe the observations and data reduction in Sec. 2, present the radial velocity monitoring in Sec. 3, update the orbital solution by combining RVs with literature astrometry in Sec. 4, and summarize our conclusions in Sec. 5.
2 Observations and data reduction
We observed the 51 Eri system at four epochs with VLT/HiRISE (see Table 3), following the same observing strategy as for AF Lep b described in Denis et al. (2025). For the first three epochs, sky backgrounds were acquired after the science exposures to subtract the leakage associated with the HiRISE MACAO guide fiber (Vigan et al. 2024). After a technical intervention in January 2025, the leakage issue was fully resolved, and subsequent observations relied on internal daytime backgrounds.
[…]
Standard CRIRES calibrations were acquired automatically the next morning based on the science observations of each night. This includes dark, flat fields and wavelength calibration files, as detailed in the CRIRES calibration plan. HiRISE observations do not require any specific internal calibrations in addition to standard daily calibrations of the spectrograph.
Data reduction was performed using the public HiRISE pipeline (Costes et al. 2024), which applies background and flat-field corrections, combines exposures, extracts the spectra, recalibrates the wavelength, and corrects for barycentric motion, while providing calibrated stellar and companion spectra22footnotemark: 2, sky and instrumental response, and associated uncertainties. The pipeline is described in detail in Denis et al. (2025) and now includes an optimal extraction scheme (Horne 1986), representing a significant improvement over earlier reductions.
3 Determination of the radial velocity of 51 Eri b
3.1 Forward modeling analysis
To measure the radial velocity (RV) of the companion, we compared the data to self-consistent atmospheric models using the forward-modeling code ForMoSA (Petrus et al. 2021). ForMoSA operates within a Bayesian framework and allows the joint analysis of spectroscopic and photometric data over a wide range of wavelengths and resolutions. Atmospheric model grids spanning effective temperature, gravity, metallicity, and related parameters are interpolated at each iteration, with optional modifications such as Doppler shifts and rotational broadening applied before computing the likelihood. The output consists of the best-fitting atmospheric parameters and their uncertainties.
We adopted the radiative–convective equilibrium model Exo-REM (Charnay et al. 2018) recomputed at a spectral resolution of 200 000, using low-resolution Exo-REM volume mixing ratio profiles in tandem with Exo_k. This model is referred to as Exo-REM/Exo_k (Radcliffe et al., in prep). A cloud-free model was used, with priors informed by Balmer et al. (2025). Given the high spectral resolution and narrow wavelength coverage, the data are expected to probe the upper atmospheric layers above the cloud base, as discussed in Xuan et al. (2022) and Denis et al. (2025). To compare the model to the data, we followed the framework described in Denis et al. (2025).
The results are summarized in Table 1, which reports the planetary RV corrected for barycentric motion and the stellar RV ($12.6\pm 0.3$$\mathrm{km}\,\mathrm{s}^{-1}$; Gontcharov 2006), as well as the planetary $v\sin i$ . The $v\sin i$ measured at the fourth epoch is inconsistent with the other epochs, likely due to instrumental fringing (see Appendix B). We therefore rely on the first three epochs to estimate the rotational broadening and derive $\hbox{$v\sin i$}=9.9\pm 2.3$$\mathrm{km}\,\mathrm{s}^{-1}$, keeping in mind that the fringing effect still affect these data. Assuming $\sin i=1$ and adopting the planetary radius $R_{p}=1.3\pm 0.03,R_{\mathrm{Jup}}$ (Balmer et al. 2025), we can derive an upper limit on the rotation period of the planet ($\sim 16.4\pm 3.8$ h). We thus estimate $P_{\mathrm{rot}}\lesssim 24$ h at 2 $\sigma$.
Table 1: Results on 51 Eri b for the 4 nights of observation333
[…]
3.2 Cross-correlation function
To assess the presence of individual molecules in the atmosphere, it is convenient to use cross-correlation function (CCF) analysis, where we correlate molecular templates with the data. In the H band, at the temperature of 51 Eri b, the dominant species are H2O, CH4 and CO. To generate the templates for these molecules, we used the pressure-temperature profiles and the volume mixing ratio profiles of Exo-REM in tandem with Exo_K. For the derivation of the CCF, we refer to Appendix C.
There appears to be no correlation between the CCF and the derived RV uncertainty. For the first and third epochs, where detection confidence is highest, the derived RV uncertainties are more than twice those of the final epoch. Although it is natural to expect RV uncertainty to correlate with CCF S/N, we are in a regime where resolution is also an important parameter. The lower vsini of the last epoch compared to the other epochs keeps the effective resolution of the template higher (see Table 1). This results in sharper variations of the log likelihood near the RV of the planet and therefore less uncertainty in its determination.
Figure 1 shows the combined CCF from the 4 epochs. Before combining the 4 CCFs, we shifted each individual CCF to the restframe of the planet to account for the variations in RV across the different epochs. In addition to the individual molecular templates, we also considered a full template including H2O + CH4 + CO. In general, the molecular detection at individual epochs are marginal, but the combined CCF exhibits a reliable detection of H2O and CH4, and a tentative detection of CO, which confirms the recent findings of Madurowicz et al. (2025).
Figure 1: Combined CCF from the 4 nights. The combined CCF has been computed by shifting the each individual CCF to a RV of 0 $\mathrm{km}\,\mathrm{s}^{-1}$ in order to account for the variations of the RV of the companion across the different epochs.
4 Updated orbital solution
Figure 2: Orbit of 51 Eri b. The left panel shows the predicted RV of the planet as a function of time. The HiRISE RV measurements are overplotted as red dots, along with the rejected orbits in grey from the addition of the RV measurements. The middle panel is a zoomed-in version of the left panel around our RV measurements. In particular, we see that a future RV measurement at the end of 2026 should enable us to reject the last low-eccentric orbital populations. This is illustrated by the dashed vertical line overplotted at the end of 2026. The right panel displays the astrometric orbit of the planet inferred, including the line of nodes and the phase of the planet. The red crosses represent the estimate of the position of the planet during the HiRISE observations.
Table 2: Orbital parameters inferred in this work.444\$a\$\$a\$footnotetext: The periastron passage is computed as $\tau=\frac{t_{\mathrm{p}}-t_{\mathrm{ref}}}{P}$ where tref is the reference time at epoch J2020.0 ($t_{\mathrm{ref}}$ = 58849 MJD) and tp is the time of periastron.
Parameter Posterior $\chi^{2}_{\mathrm{min}}$
Semi-major axis $a$ (au) $9.08_{-0.19}^{+0.57}$ 8.91
Eccentricity $e$ $0.55_{-0.07}^{+0.03}$ 0.58
Inclination $i$ (°) $159.24_{-11.48}^{+5.74}$ 165
Argument of periastron $\omega$ (°) $55.41_{-30.75}^{+51.00}$ 84
Ascending node $\Omega$ (°) $58.25_{-47.60}^{+80.80}$ 69
period P (years) $3.49_{-0.86}^{+1.74} 22$.76
[…]
To update the orbit of 51 Eri b, we compiled relative astrometric data from Gemini/GPI and Keck/NIRC2 (Macintosh et al. 2015; Rosa et al. 2020), VLT/SPHERE (Maire et al. 2019) and JWST/NIRCam (Balmer et al. 2025). We used the Markov Chain Monte Carlo (MCMC) framework designed by Beust et al. (2016) that makes use of universal Keplerian variables with Stumpff functions (Danby 1987). This formulation achieves more efficient convergence in the case of astrometric orbits sampled over a small orbital arc, as in the present case. The MCMC was run using 8 walkers. The burn-in and convergence was assessed using Gelman-Rubin statistics (Gelman et al. 2003; Ford 2006). Before beginning the inference of the chains, we require that $\hat{R}(z)<1.1$ and $\hat{T}(z)>100$ (see Sects. 3.3 and 3.4 of Ford 2006). Then, we let the chains be filled until $\hat{R}(z)<1.01$ and $\hat{T}(z)>1000$. The filled chains constitutes the output of the orbital fit.
We do not directly fit for classical orbital elements: the semi-major axis $a$, the eccentricity $e$, the inclination $i$, the argument of periastron $\omega$, the longitude of ascending node $\Omega$, and the time of periastron passage $t_{p}$. Ford (2006) suggests that working with combinations of these elements allows an optimization of the Markov Chains. We specifically fit for $\ln q$, $e\cos(\Omega+\omega)$, $e\sin(\Omega+\omega)$, $sin(i/2)\cos\Omega$, $\sin(i/2)\sin\Omega$, $\ln P$, and $s$. Here $q$ is the periastron, $P_{q}$ is the orbital period of a circular orbit with radius $q$, and $s$ is the universal variable (a reformulation of the eccentric anomaly, Danby 1987) at a reference epoch. Note that $P$ can also be defined as $P_{q}=2\pi qv_{q}$, where $v_{q}$ is the velocity at periastron. From any solution with those fitted parameters, the classical elements can be derived unambiguously.
The updated orbit and associated parameters are presented in Fig. 2 and Table 2. The full posterior distributions of orbital parameters are presented in Fig. 6. The left panel of Fig. 2 shows the predicted RV differences between the planet and the star as a function of time. The middle panel is a zoomed-in version centered on the HiRISE measurements and the right panel displays the astrometric orbit of the companion. These plots are generated using 500 orbits randomly drawn from the posterior distributions.
The inclusion of our RV measurements rules out a significant number of orbital solutions, as can be seen on the left panel of Fig. 2. The middle panel shows that a future RV measurement of the planet at the end of 2026 will potentially completely reject the last population of low-eccentricity orbits and place even tighter constraints on the eccentricity. As illustrated in Fig. 6, adding the HiRISE RV measurements significantly improves the orbit. We now estimate the eccentricity to lie within the range 0.2–0.7 at $5\sigma$, with a median at $e=0.55$, completely ruling out very eccentric orbits, above 0.7.
The HiRISE measurements also allow to break the typical degeneracy arising with the argument of periastron, $\omega$, and the longitude of ascending node, $\Omega$, when using only astrometric data to fit orbits of exoplanets. This is because the computation of the astrometry of the planet from the orbital parameters is invariant to the transformation ($\omega^{\prime}=\omega+\pi$, $\Omega^{\prime}=\Omega+\pi$), typically producing double-peaked structures in the posterior distributions of these parameters. As a consequence, including our RV measurements enable the determination of the direction of the motion of the companion along the line of sight. This allows us to compute a phase curve of the companion (see the right panel of Fig. 7), which is essential for planning follow-up observations using reflected-light imaging or spectroscopy with future instrumentation. This is particularly pertinent in this case, since the reflected-light contrast is predicted to be higher than the thermal-light contrast for this planet, mostly under favourable phase functions (see Appendix. E).
5 Discussion and conclusions
[…]
The implications of the planet’s high eccentricity have been extensively discussed in previous studies (Maire et al. 2019; Dupuy et al. 2021; Balmer et al. 2025). Dynamical interactions with a massive inner companion could in principle reproduce the observed eccentricity, but the required mass would exceed SPHERE detection limits (Maire et al. 2019). An alternative scenario involves secular resonances with the massive circumstellar disk, which can excite planetary eccentricities up to $\sim$0.6 even for modest disk eccentricities (Li and Lai 2023). Kozai–Lidov oscillations induced by a distant third body, such as the stellar companion GJ 3305 AB, can also generate large eccentricities (Kozai 1962; Lidov 1962). However, the wide separation of GJ 3305 AB implies a Kozai–Lidov timescale longer than the age of the system, making this scenario unlikely (Montet et al. 2015), except in the case of a very high stellar companion eccentricity ($\gtrsim 0.9$; see Appendix F). Kozai–Lidov oscillations with an undetected low-mass outer companion remain possible; given SPHERE/IRDIS detection limits (Chomez et al. 2025), such a companion would likely be Jovian or sub-Jovian ($\lesssim 1$$2,M_{\mathrm{Jup}}$). Continued astrometric and radial velocity monitoring will further constrain this scenario (Lacour et al. 2021).
If Kozai–Lidov oscillations can be ruled out, past dynamical interactions, such as a stellar flyby (Pfalzner et al. 2024), offer another possible explanation. Further N-body simulations and searches for potential stellar perturbers in Gaia data, as performed for the HD 106906 system (De Rosa and Kalas 2019), will provide additional insight into the formation and dynamical evolution of the 51 Eri system.
Acknowledgements.
This project has received funding from Agence Nationale de la Recherche (ANR) under grant ANR-23-CE31-0006-01 (MIRAGES). The HiRISE instrument has been developed with funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreements No. 757561 (HiRISE) and 678777 (ICARUS), from the Commission Spécialisée Astronomie-Astrophysique (CSAA) of CNRS/INSU, and from the Action Spécifique Haute Résolution Angulaire (ASHRA) of CNRS/INSU co-funded by CNES, and from Région Provence-Alpes-Côte d’Azur under grant agreement 2014-0276 (ASOREX). This research has made use of computing facilities operated by CeSAM data center at LAM, Marseille, France. S. Petrus was supported by an appointment to the NASA Postdoctoral Program at the NASA-Goddard Space Flight Center, administered by Oak Ridge Associated Universities under contract with NASA. N. Godoy acknowledges funding from the European Union (ERC, ESCAPE, project No 101044152).
Appendix A Observations of 51 Eri b
A summary of the HiRISE observations of 51 Eri b is presented in Table 3.
Table 3: 51 Eri observations.555\$a\$\$a\$footnotetext: The transmission is measured on the star. The value provided corresponds to the 95th-percentile of the transmission computed over the whole spectral range covered with HiRISE.\$b\$\$b\$footnotetext: The angular separation was computed from the orbital fitting (see Sect. 4).
UT date Setting Integration time Integration time Seeing Transmission\$a\$\$a\$footnotetext: Angular separation\$b\$\$b\$footnotetext: on star on companion
[min] [min] [′′ ] [%] [mas]
2023-11-21 H1567 4 100 0.6 4.0 295
2024-12-01 H1567 4 120 0.6 - 0.7 3.9 244
2025-02-03 H1567 12 120 0.5 - 0.9 3.9 234
2025-09-11 H1567 4 120 0.5 - 0.8 3.7 200
Appendix B Considering whether the results of the fourth epoch are biased
In Table 1, we see that the $v\sin i$ of the last epoch is particularly inconsistent with the other epochs. A possible explanation of this bias is the presence of spectral fringes in the data due to the dichroic filter present in HiRISE (see Fig. 1 of El Morsy et al. 2022). This dichroic filter is a plane-parallel piece of glass with anti-reflection coatings, which has the unfortunate property of acting as a Fabry-Pérot cavity and causing oscillations as a function of wavelength in the stellar speckles signal. This is identified in the power-spectral density (PSD) of the data filtered from the stellar speckle estimates, which is plotted for our four epochs in Fig. 3. The peaks located at a spectral resolution of $\sim$25 000 correspond to the fringes. The equivalent wavelength in the velocity space is $\approx$12 $\mathrm{km}\,\mathrm{s}^{-1}$. It means that the fringes induce a periodic signal in the spectrum whose peaks are separated by $\sim$12 $\mathrm{km}\,\mathrm{s}^{-1}$ in the velocity dimension. This translates into correlated structures in the CCF at multiples of $\sim$12 $\mathrm{km}\,\mathrm{s}^{-1}$.
To highlight these correlated structures, it is convenient to compute the Autocorrelation Function (ACF) of the CCF. Fig. 4 presents the individual ACFs of the 4 CCFs. The first three epochs are very consistent. The last epoch, however, exhibits prominent correlated structures at velocities multiples of $\sim$ 11 $\mathrm{km}\,\mathrm{s}^{-1}$, which is consistent with the previous analysis. It also shows that the fringes cause a decrease in the Full Width at Half-Maximum (FWHM) of the ACF, compared to the other epochs. It means that the fringes bias the FWHM of the CCF to lower values. Since we expect the FWHM of the CCF and the $v\sin i$ to be correlated (see Brady et al. 2023, for example), we can conclude that the fringes likely bias the $v\sin i$ estimate of the last epoch towards lower values. However, they likely do not bias the RV as much. Provided that the template correlates better with the data than the fringes do, the CCF still provides a correct measurement of the RV of the planet.
This effect is actually quite common and has been identified in other contexts such as in the KPIC instrument (Horstman et al. 2024) or in the JWST/MIRI medium-resolution spectroscopy mode (Martos et al. 2025). This systematic noise can become a significant limitation in the detection of the planet and its intensity is proportional to the intensity of the stellar light in the speckle field at the location of the companion. If we compute the angular separation of the companion at each of our observing epochs (see Table 3), we see that the planet was at its closest to the host star during the last epoch. And indeed looking at the PSD in Fig. 3, this same epoch is also the one where the peak corresponding to the fringes is the strongest.
A potential solution would to be include a model of the fringing signal, as in Horstman et al. (2024). However, this is beyond the scope of the present paper and we consider the $v\sin i$ estimate of the fourth epoch to be unreliable.
Figure 3: Power-spectral density (PSD) of the data filtered from the stellar speckles estimate of our four epochs. The PSD was lowpass filtered in order to focus on its low-frequency content. A secondary axis is displayed at the top, which corresponds to an equivalent velocity. A peak is visible around a resolution of 25 000 for each epoch, which corresponds to $\sim$12 $\mathrm{km}\,\mathrm{s}^{-1}$. This peak is particularly prominent in the data from the last epoch, when the planet was closer to the host star.
Figure 4: Normalized Autocorrelation Functions (ACF) of the individual CCFs. The first three epochs show highly consistent ACFs, whereas the last epoch shows deeply correlated structures in the CCF at multiples of 11 km/s. The vertical dashed lines are plotted at multiples of $\pm$ 11 $\mathrm{km}\,\mathrm{s}^{-1}$ to highlight the correlated structures.
Appendix C Method for computing the cross-correlation functions
To compute the cross-correlation functions, we use the formalism defined in Zucker (2003). The cross-correlation function is defined as:
$CCF(RV)=\frac{\sum_{i}d_{HF}(i)t_{HF}^{RV}(i)}{\sqrt{\sum_{i}d_{HF}(i)^{2}}\sqrt{\sum_{i}t_{HF}^{RV}(i)^{2}}}.$ (1)
[…]
Figure 5: Individual CCF for the 4 epochs.
Appendix D Orbit posterior distribution
Fig. 6 presents the posterior distribution of the orbital parameters of 51 Eri b, without our RV measurements (in red) and with our RV measurements (in blue). The inclusion of the HiRISE RV measurements brings significant improvements to the orbital parameters. It also breaks the degeneracy between particular values of $\omega$ and $\Omega$ typically arising with orbital fittings using only astrometric data points.
Figure 6: Posterior distribution of the inferred parameters of 51 Eri b. The posterior distribution in red was obtained including all available relative astrometry measurements. The posterior distribution in blue includes the HiRISE RV measurements.
Appendix E Contrast in reflected and emitted light
In order to determine whether the planet could be a target of interest for reflected-light imaging and spectroscopy, it is crucial to establish whether the reflected-light contrast overpowers the thermal light contrast. We compared the predicted reflected-light contrast with the thermal light contrast within the visible spectrum. The reflected-light contrast was computed using the following formula:
$C=f(\phi)\,A\,\left(\frac{R}{a}\right)^{2},$ (2)
with $R$ the radius of the planet, $a$ its semi-major axis, $A$ its albedo and $f$ the phase function. We considered 3 different values for the phase function to show the impact of the phase of the planet. For $A$, we considered a value of 0.5, similar to that of Jupiter in visible (Karkoschka 1998). The values of $a$ and $R$ have been determined from the orbital and atmospheric fit, respectively.
To compute the thermal light contrast in visible of the planet, we first compared the low-resolution spectro-photometric observations of 51 Eri b with the self-consistent model Exo-REM in ForMoSA. We used the Y/J/H1-band band spectra from VLT/SPHERE (Brown-Sevilla et al. 2023), the Keck/GPI K1 and K1-bands spectroscopy and the Keck/NIRC2 L and M-bands photometry (Rajan et al. 2017) and the JWST/NIRCam F410M photometry (Balmer et al. 2025). The results are consistent with the ones presented in Balmer et al. (2025). Then, because the Exo-REM model is not extended in the visible light range, we computed the flux in the visible light range using petitRADTRANS (Mollière et al. 2019) with the best Exo-REM PT profile and volume mixing ratio inferred from the forward modeling.
The results are presented in Fig. 7 (right). For favourable phase angles of the planet, we see that the predicted reflected-light contrast is above the thermal light contrast. For very unfavourable phase angles of the planet, there is some regions where the thermal light is above the predicted reflected light, but between 0.5 and 0.7 $\upmu\mathrm{m}$, the reflected-light contrast is still predominant compared to the thermal light contrast. These conclusions do not dramatically change even when considering a very low albedo, such as 0.1.
[…]
Figure 8: Kozai–Lidov oscillation timescale as a function of the perturber mass and semi-major axis for different eccentricities. The SPHERE/IRDIS detection limits are shown in white; perturbers located to the right of this curve (white hatched region) would be detectable by SPHERE. The red curve indicates the age of the system, with the red hatched region corresponding to Kozai–Lidov timescales longer than the system age. The position of the stellar binary companion GJ 3305 AB is also indicated.
Appendix F Kozai-Lidov oscillations with 51 Eri b
In the Kozai-Lidov scenario, the orbit of a planet is perturbed by a distant third body, which can lead to oscillations in the inclination and eccentricity of the inner companion. The timescale of such oscillations is on the order of magnitude of (Ford and Rasio 2008):
$P_{\mathrm{Kozai}}\approx P_{\mathrm{planet}}\frac{M_{\star}}{M_{\mathrm{pert}}}\left(\frac{a_{\mathrm{pert}}}{a_{\mathrm{planet}}}\right)^{3}\left(1-e_{\mathrm{pert}}^{2}\right)^{3/2}.$ (3)
In this equation, $P_{\mathrm{Kozai}}$ is the period of the Kozai-Lidov oscillations, $M_{\star}$, the mass of the host star, $M_{\mathrm{pert}}$, $a_{\mathrm{pert}}$ and $e_{\mathrm{pert}}$ the mass, semi-major axis and eccentricity of the distant perturber, and $P_{\mathrm{planet}}$ and $a_{\mathrm{planet}}$ the period and semi-major axis of the inner planet. We computed the period of the Kozai-Lidov oscillations for different values of $M_{\mathrm{pert}}$, $a_{\mathrm{pert}}$ and $e_{\mathrm{pert}}$ using the posterior distributions of $P_{\mathrm{planet}}$ and $a_{\mathrm{planet}}$ inferred in this study. Fig. 8 shows the results of this analysis. The red curve shows the age of the system. The parameter space for Kozai-Lidov oscillations on a timescale longer than the age of the system lies above this curve. Therefore, this parameter space is not applicable. The white curve represents the SPHERE detection limits from Chomez et al. (2025). Hence, the range of possible solutions for the mass and semi-major axis of the unseen perturber lies between these 2 curves. The binary stellar companion GJ 3305 AB is overplotted. We considered the instantaneous sky-projected separation as the semi-major axis of the system and ($\approx$ 2000 AU, Feigelson et al. 2006) and the dynamical mass estimate of the binary ($\approx$ 1.1 M, Montet et al. 2015). If the binary stellar companion has a high eccentricity ($\gtrsim$ 0.9), it could potentially be responsible for the orbit of 51 Eri b, especially since, with such a high eccentricity, the semi-major axis could be significantly lower than the value of 2000 AU used in this analysis, lowering the Kozai-Lidov period. Indeed, the maximum possible separation of the stellar companion is $r_{apastron}=(1+e)*a$. This provides a lower limit for the semi-major axis of the stellar companion, i.e. $a>\frac{r_{apastron}}{1+e}$. For extreme values of the eccentricity, the minimum possible semi-major axis is $a_{min}=\frac{r_{apastron}}{2}$. This yields a range of possible values for the semi-major axis of the stellar companion, which can be as low as half of the value of the instantaneous sky-projected separation. This is a very unlikely scenario, however. Additional constrains on the possibility of GJ 3305 AB being in Kozai-Lidov resonance with 51 Eri b could be brought by a direct measurement of the RV of the stellar companion, as it would help constrain both the inclination and the longitude of ascending node of the stellar companion. The Kozai-Lidov resonance requires the mutual inclination I between the 2 objects in resonance satisfies (Kozai 1962; Lidov 1962):
$39\degree<I<141\degree.$ (4)
From the RV measurement of the stellar companion, we can better constrain its inclination and longitude of ascending node. Then, the mutual inclination between 51 Eri b and GJ 3305 AB is computed as (Murray and Dermott 2000):
$\cos I=\cos i_{1}\cos i_{2}+\sin i_{1}\sin i_{2}\cos(\Omega_{1}-\Omega_{2}),$ (5)
with i1, i2, $\Omega_{1}$, $\Omega_{2}$ the inclinations and longitude of ascending node of the 2 companions.

### v2

@@
-Eccentric and cool: a high-spectral resolution view of
+Eccentric and cool: A high-spectral-resolution view of
 11institutetext: Aix Marseille Univ, CNRS, CNES, LAM, Marseille, France
@@
 2, 85748 Garching, Germany 2222institutetext: Physics & Astronomy Dpt, University of Exeter, Exeter, EX4 4QL, UK 2323institutetext: Institute for Astronomy, University of Hawaii at Manoa, Honolulu, HI 96822, USA
-Eccentric and cool: a high-spectral resolution view of
+Eccentric and cool: A high-spectral-resolution view of
 51 Eri b with VLT/HiRISE [^1] Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programmes ID 112.25FU, 114.2712, and 115.284P
@@
 Zins
-(Received 05 January 2026; accepted 31 January 2026)
-Discovered almost 10 years ago, the giant planet 51 Eridani b is one of the least separated ($\approx$0.2′′) and faintest (J $\approx$ 19.74 mag) directly imaged exoplanets known to date.
-Its atmospheric properties have been thoroughly investigated through low- and medium-resolution spectroscopic observations, enabling robust characterization of the planet’s bulk parameters.
+(Received 05 January 2025; accepted 31 January 2026)
+Discovered almost ten years ago, the giant planet 51 Eridani b is one of the least separated ($\approx$0.2′′) and faintest (J $\approx$ 19.74 mag) directly imaged exoplanets known to date.
+Its atmospheric properties have been thoroughly investigated through low- and medium-resolution spectroscopic observations, enabling the robust characterization of the planet’s bulk parameters.
 However, the planet’s intrinsically high contrast renders high-resolution spectroscopic observations difficult, despite their potential to yield key measurements essential for a more comprehensive characterization.
-This study seeks to constrain the planet’s radial velocity, enabling a full 3D orbital solution when integrated with previous measurements.
-We have obtained 4 high-contrast high-resolution (R $\approx$ 140 000) spectroscopic datasets of the planet, collected over a two-year interval with the HiRISE visitor instrument at the VLT to derive the planet’s radial velocity.
-Using self-consistent models of atmosphere, we were able to derive the radial velocity of the planet at each of the 4 epochs.
-These radial velocity measurements were then used in combination with all existing relative astrometry in order to constrain the orbit of the planet.Our radial velocity measurements allow us to break the degeneracy along the line of sight, making it now possible the unambiguous interpretation of the phase curve of the companion.
-We further constrain the orbital parameters, particularly the eccentricity, for which we derive e = $0.55_{-0.07}^{+0.03}$.
+This study sought to constrain the planet’s radial velocity, enabling a full 3D orbital solution when integrated with previous measurements.
+We obtained four high-contrast, high-resolution (R $\approx$ 140 000) spectroscopic datasets of the planet, collected over a two-year interval with the HiRISE visitor instrument at the VLT to derive the planet’s radial velocity.
+Using self-consistent models of atmosphere, we were able to derive the radial velocity of the planet at each of the four epochs.
+These radial-velocity measurements were then used in combination with all existing relative astrometry in order to constrain the orbit of the planet.Our radial velocity measurements allowed us to break the degeneracy along the line of sight, making the unambiguous interpretation of the phase curve of the companion possible.
+We further constrained the orbital parameters, and particularly the eccentricity, for which we derive e = $0.55_{-0.07}^{+0.03}$.
 The relatively high eccentricity indicates that the system has experienced dynamical interactions induced by an external perturber.
-We place constraints on the mass and semi-major axis of a hypothetical, unseen outer planet capable of producing the observed high eccentricities.
+We place constraints on the mass and semimajor axis of a hypothetical, unseen outer planet capable of producing the observed high eccentricities.
 Key Words.:
-Instrumentation: high angular resolution – Instrumentation: adaptive optics – Instrumentation: spectrographs – Techniques: high-angular resolution – Techniques: spectroscopy – Infrared: planetary systems
+Instrumentation: high angular resolution – Instrumentation: adaptive optics – Instrumentation: spectrographs – Techniques: high angular resolution – Techniques: spectroscopy – Infrared: planetary systems
 1 Introduction
@@
 Eccentricity, for instance, is an efficient indicator of planetary formation and dynamical evolution.
-While planets forming in protoplanetary disks are expected to form on circular orbits, they can develop non-zero eccentricities through dynamical processes such as planet–planet scattering (Ford and Rasio 2008), Kozai–Lidov oscillations (Kozai 1962; Lidov 1962), planet–disk interactions (Li and Lai 2023), or stellar flybys (Pfalzner et al.
+While planets forming in protoplanetary disks are expected to form on circular orbits, they can develop nonzero eccentricities through dynamical processes such as planet–planet scattering (Ford and Rasio 2008), Kozai–Lidov oscillations (Kozai 1962; Lidov 1962), planet–disk interactions (Li and Lai 2023), or stellar flybys (Pfalzner et al.
 2024).
-A population-level dichotomy exists between brown dwarf and giant planet companions, with brown dwarfs peaking at high eccentricities ($e\approx 0.6$$0.9$) and giant planets typically exhibiting low eccentricities (Bowler et al.
+A population-level dichotomy exists between brown-dwarf and giant-planet companions, with brown dwarfs peaking at high eccentricities ($e\approx 0.6$$0.9$) and giant planets typically exhibiting low eccentricities (Bowler et al.
 2020).
-The giant planet 51 Eridani b (hereafter 51 Eri b) is an exception, as recent studies suggest a potentially high eccentricity (Rosa et al.
+The giant planet 51 Eridani b (hereafter 51 Eri b) is an exception, as recent studies suggested a potentially high eccentricity (Rosa et al.
 2020; Dupuy et al.
 2021; Balmer et al.
-2025), indicative of past dynamical interactions.
+2025), which is indicative of past dynamical interactions.
 Discovered in 2014 by the Gemini Planet Imager Exoplanet Survey (Macintosh et al.
@@
 2024).
-The system also hosts the wide, co-moving M-dwarf binary GJ 3305 AB at a projected separation of $\sim$2000,au (Feigelson et al.
+The system also hosts the wide, co-moving M-dwarf binary GJ 3305 AB at a projected separation of $\sim$2000 au (Feigelson et al.
 2006; Kasper et al.
@@
 […]
-In this work, we present nearly two years of high-resolution ($R=140,000$) direct spectroscopic observations of 51 Eri b obtained with the HiRISE instrument (Vigan et al.
-2024), combining SPHERE (Beuzit et al.
+In this paper, we present nearly two years of high-resolution ($R=140,000$) direct spectroscopic observations of 51 Eri b obtained with the HiRISE instrument (Vigan et al.
+2024), which combines SPHERE (Beuzit et al.
 2019) and CRIRES (Dorn et al.
 2023).
-We describe the observations and data reduction in Sec.
-2, present the radial velocity monitoring in Sec.
-3, update the orbital solution by combining RVs with literature astrometry in Sec.
-4, and summarize our conclusions in Sec.
+We describe the observations and data reduction in Sect.
+2 and present the radial-velocity monitoring in Sect.
+3.
+In Sect.
+4, we provide an update of the orbital solution realized by combining RVs with literature astrometry.
+We summarize our conclusions in Sect.
 5.
 2 Observations and data reduction
-We observed the 51 Eri system at four epochs with VLT/HiRISE (see Table 3), following the same observing strategy as for AF Lep b described in Denis et al.
+We observed the 51 Eri system at four epochs with VLT/HiRISE (see Table 3) following the same observing strategy for AF Lep b described in Denis et al.
 (2025).
@@
 Standard CRIRES calibrations were acquired automatically the next morning based on the science observations of each night.
-This includes dark, flat fields and wavelength calibration files, as detailed in the CRIRES calibration plan.
+These include dark, flat fields and wavelength-calibration files, as detailed in the CRIRES calibration plan.
 HiRISE observations do not require any specific internal calibrations in addition to standard daily calibrations of the spectrograph.
 Data reduction was performed using the public HiRISE pipeline (Costes et al.
-2024), which applies background and flat-field corrections, combines exposures, extracts the spectra, recalibrates the wavelength, and corrects for barycentric motion, while providing calibrated stellar and companion spectra22footnotemark: 2, sky and instrumental response, and associated uncertainties.
+2024), which applies background and flat-field corrections, combines exposures, extracts the spectra, recalibrates the wavelength, and corrects for barycentric motion, while providing calibrated stellar and companion spectra,22footnotemark: 2sky and instrumental response, and associated uncertainties.
 The pipeline is described in detail in Denis et al.
@@
 ForMoSA operates within a Bayesian framework and allows the joint analysis of spectroscopic and photometric data over a wide range of wavelengths and resolutions.
-Atmospheric model grids spanning effective temperature, gravity, metallicity, and related parameters are interpolated at each iteration, with optional modifications such as Doppler shifts and rotational broadening applied before computing the likelihood.
+Atmospheric model grids spanning effective temperature, gravity, metallicity, and related parameters are interpolated at each iteration, with optional modifications such as Doppler shifts and rotational broadening to be applied before computing the likelihood.
 The output consists of the best-fitting atmospheric parameters and their uncertainties.
@@
 2018) recomputed at a spectral resolution of 200 000, using low-resolution Exo-REM volume mixing ratio profiles in tandem with Exo_k.
-This model is referred to as Exo-REM/Exo_k (Radcliffe et al., in prep).
-A cloud-free model was used, with priors informed by Balmer et al.
+This model is referred to as Exo-REM k26 (Radcliffe et al., submitted).
+A cloud-free model was used, along with the priors of Balmer et al.
 (2025).
@@
 (2025).
-The results are summarized in Table 1, which reports the planetary RV corrected for barycentric motion and the stellar RV ($12.6\pm 0.3$$\mathrm{km}\,\mathrm{s}^{-1}$; Gontcharov 2006), as well as the planetary $v\sin i$ .
+The results are summarized in Table 1, which reports the planetary RV corrected for barycentric motion, the stellar RV ($12.6\pm 0.3$$\mathrm{km}\,\mathrm{s}^{-1}$; Gontcharov 2006), and the planetary $v\sin i$.
 The $v\sin i$ measured at the fourth epoch is inconsistent with the other epochs, likely due to instrumental fringing (see Appendix B).
-We therefore rely on the first three epochs to estimate the rotational broadening and derive $\hbox{$v\sin i$}=9.9\pm 2.3$$\mathrm{km}\,\mathrm{s}^{-1}$, keeping in mind that the fringing effect still affect these data.
+We therefore relied on the first three epochs to estimate the rotational broadening and derive $\hbox{$v\sin i$}=9.9\pm 2.3$$\mathrm{km}\,\mathrm{s}^{-1}$, keeping in mind that the fringing effect still affects these data.
 Assuming $\sin i=1$ and adopting the planetary radius $R_{p}=1.3\pm 0.03,R_{\mathrm{Jup}}$ (Balmer et al.
 2025), we can derive an upper limit on the rotation period of the planet ($\sim 16.4\pm 3.8$ h).
-We thus estimate $P_{\mathrm{rot}}\lesssim 24$ h at 2 $\sigma$.
-Table 1: Results on 51 Eri b for the 4 nights of observation333
+We thus estimated $P_{\mathrm{rot}}\lesssim 24$ h at 2 $\sigma$.
+Table 1: Results on 51 Eri b for the our nights of observations.333
 […]
@@
 3.2 Cross-correlation function
-To assess the presence of individual molecules in the atmosphere, it is convenient to use cross-correlation function (CCF) analysis, where we correlate molecular templates with the data.
-In the H band, at the temperature of 51 Eri b, the dominant species are H2O, CH4 and CO.
-To generate the templates for these molecules, we used the pressure-temperature profiles and the volume mixing ratio profiles of Exo-REM in tandem with Exo_K.
-For the derivation of the CCF, we refer to Appendix C.
+To assess the presence of individual molecules in the atmosphere, it is convenient to use a cross-correlation-function (CCF) analysis, in which we correlate molecular templates with the data.
+In the H band, at the temperature of 51 Eri b, the dominant species are H2O, CH4, and CO.
+To generate the templates for these molecules, we used the pressure-temperature profiles and the volume-mixing ratio profiles of Exo-REM in tandem with Exo_K.
+For the derivation of the CCF, we refer the reader to Appendix C.
 There appears to be no correlation between the CCF and the derived RV uncertainty.
 For the first and third epochs, where detection confidence is highest, the derived RV uncertainties are more than twice those of the final epoch.
-Although it is natural to expect RV uncertainty to correlate with CCF S/N, we are in a regime where resolution is also an important parameter.
+Although it is natural to expect RV uncertainty to correlate with the CCF Signal-to-Noise (S/N), we are in a regime where resolution is also an important parameter.
 The lower vsini of the last epoch compared to the other epochs keeps the effective resolution of the template higher (see Table 1).
 This results in sharper variations of the log likelihood near the RV of the planet and therefore less uncertainty in its determination.
-Figure 1 shows the combined CCF from the 4 epochs.
-Before combining the 4 CCFs, we shifted each individual CCF to the restframe of the planet to account for the variations in RV across the different epochs.
+Figure 1 shows the combined CCF from the four epochs.
+Before combining the four CCFs, we shifted each individual CCF to the rest frame of the planet to account for the variations in RV across the different epochs.
 In addition to the individual molecular templates, we also considered a full template including H2O + CH4 + CO.
-In general, the molecular detection at individual epochs are marginal, but the combined CCF exhibits a reliable detection of H2O and CH4, and a tentative detection of CO, which confirms the recent findings of Madurowicz et al.
+In general, the molecular detections at individual epochs are marginal, but the combined CCF exhibits a reliable detection of H2O and CH4 and a tentative detection of CO, which confirms the recent findings of Madurowicz et al.
 (2025).
-Figure 1: Combined CCF from the 4 nights.
-The combined CCF has been computed by shifting the each individual CCF to a RV of 0 $\mathrm{km}\,\mathrm{s}^{-1}$ in order to account for the variations of the RV of the companion across the different epochs.
+Figure 1: Combined CCF from the four nights.
+The combined CCF was computed by shifting the each individual CCF to an RV of 0 $\mathrm{km}\,\mathrm{s}^{-1}$ in order to account for the variations of the RV of the companion across the different epochs.
 4 Updated orbital solution
@@
 The red crosses represent the estimate of the position of the planet during the HiRISE observations.
-Table 2: Orbital parameters inferred in this work.444\$a\$\$a\$footnotetext: The periastron passage is computed as $\tau=\frac{t_{\mathrm{p}}-t_{\mathrm{ref}}}{P}$ where tref is the reference time at epoch J2020.0 ($t_{\mathrm{ref}}$ = 58849 MJD) and tp is the time of periastron.
+Table 2: Orbital parameters inferred in this work.444\$a\$\$a\$footnotetext: The periastron passage is computed as $\tau=\frac{t_{\mathrm{p}}-t_{\mathrm{ref}}}{P}$, where tref is the reference time at epoch J2020.0 ($t_{\mathrm{ref}}$ = 58849 MJD) and tp is the time of periastron.
 Parameter Posterior $\chi^{2}_{\mathrm{min}}$
@@
 2020), VLT/SPHERE (Maire et al.
-2019) and JWST/NIRCam (Balmer et al.
+2019), and JWST/NIRCam (Balmer et al.
 2025).
 We used the Markov chain Monte Carlo (MCMC) framework designed by Beust et al.
-(2016) that makes use of universal Keplerian variables with Stumpff functions (Danby 1987).
+(2016), which makes use of universal Keplerian variables with Stumpff functions (Danby 1987).
 This formulation achieves more efficient convergence in the case of astrometric orbits sampled over a small orbital arc, as in the present case.
-The MCMC was run using 8 walkers.
+The MCMC was run using eight walkers.
 The burn-in and convergence was assessed using Gelman-Rubin statistics (Gelman et al.
 2003; Ford 2006).
-Before beginning the inference of the chains, we require that $\hat{R}(z)<1.1$ and $\hat{T}(z)>100$ (see Sects.
+Before beginning the inference of the chains, we required that $\hat{R}(z)<1.1$ and $\hat{T}(z)>100$ (see Sects.
 3.3 and 3.4 of Ford 2006).
 Then, we let the chains be filled until $\hat{R}(z)<1.01$ and $\hat{T}(z)>1000$.
-The filled chains constitutes the output of the orbital fit.
-We do not directly fit for classical orbital elements: the semi-major axis $a$, the eccentricity $e$, the inclination $i$, the argument of periastron $\omega$, the longitude of ascending node $\Omega$, and the time of periastron passage $t_{p}$.
-Ford (2006) suggests that working with combinations of these elements allows an optimization of the Markov Chains.
+The filled chains constitute the output of the orbital fit.
+We did not directly fit for classical orbital elements: the semimajor axis, $a$; the eccentricity, $e$; the inclination, $i$; the argument of periastron, $\omega$; the longitude of the ascending node, $\Omega$; and the time of periastron passage, $t_{p}$.
+Ford (2006) suggested that working with combinations of these elements allows an optimization of the Markov chains.
 We specifically fit for $\ln q$, $e\cos(\Omega+\omega)$, $e\sin(\Omega+\omega)$, $sin(i/2)\cos\Omega$, $\sin(i/2)\sin\Omega$, $\ln P$, and $s$.
-Here $q$ is the periastron, $P_{q}$ is the orbital period of a circular orbit with radius $q$, and $s$ is the universal variable (a reformulation of the eccentric anomaly, Danby 1987) at a reference epoch.
+Here, $q$ is the periastron, $P_{q}$ is the orbital period of a circular orbit with radius $q$, and $s$ is the universal variable (a reformulation of the eccentric anomaly; Danby 1987) at a reference epoch.
 Note that $P$ can also be defined as $P_{q}=2\pi qv_{q}$, where $v_{q}$ is the velocity at periastron.
-From any solution with those fitted parameters, the classical elements can be derived unambiguously.
+From any solution with those fit parameters, the classical elements can be derived unambiguously.
 The updated orbit and associated parameters are presented in Fig.
@@
 2 shows the predicted RV differences between the planet and the star as a function of time.
-The middle panel is a zoomed-in version centered on the HiRISE measurements and the right panel displays the astrometric orbit of the companion.
-These plots are generated using 500 orbits randomly drawn from the posterior distributions.
-The inclusion of our RV measurements rules out a significant number of orbital solutions, as can be seen on the left panel of Fig.
+The middle panel shows a zoomed-in view centered on the HiRISE measurements, and the right panel displays the astrometric orbit of the companion.
+These plots were generated using 500 orbits randomly drawn from the posterior distributions.
+The inclusion of our RV measurements rules out a significant number of orbital solutions, as can be seen in the left panel of Fig.
 2.
@@
 6, adding the HiRISE RV measurements significantly improves the orbit.
-We now estimate the eccentricity to lie within the range 0.2–0.7 at $5\sigma$, with a median at $e=0.55$, completely ruling out very eccentric orbits, above 0.7.
-The HiRISE measurements also allow to break the typical degeneracy arising with the argument of periastron, $\omega$, and the longitude of ascending node, $\Omega$, when using only astrometric data to fit orbits of exoplanets.
+We now estimate the eccentricity to lie within the 0.2–0.7 range at $5\sigma$, with a median at $e=0.55$, completely ruling out very eccentric orbits, above 0.7.
+The HiRISE measurements also allowed us to break the typical degeneracy arising with the argument of periastron, $\omega$, and the longitude of ascending node, $\Omega$, when using only astrometric data to fit orbits of exoplanets.
 This is because the computation of the astrometry of the planet from the orbital parameters is invariant to the transformation ($\omega^{\prime}=\omega+\pi$, $\Omega^{\prime}=\Omega+\pi$), typically producing double-peaked structures in the posterior distributions of these parameters.
-As a consequence, including our RV measurements enable the determination of the direction of the motion of the companion along the line of sight.
-This allows us to compute a phase curve of the companion (see the right panel of Fig.
+As a consequence, including our RV measurements enabled the determination of the direction of the motion of the companion along the line of sight.
+This allowed us to compute a phase curve of the companion (see the right panel of Fig.
 7), which is essential for planning follow-up observations using reflected-light imaging or spectroscopy with future instrumentation.
-This is particularly pertinent in this case, since the reflected-light contrast is predicted to be higher than the thermal-light contrast for this planet, mostly under favourable phase functions (see Appendix.
+This is particularly pertinent in this case, since the reflected-light contrast is predicted to be higher than the thermal-light contrast for this planet, mostly under favorable phase functions (see Appendix.
 E).
@@
 2019).
-An alternative scenario involves secular resonances with the massive circumstellar disk, which can excite planetary eccentricities up to $\sim$0.6 even for modest disk eccentricities (Li and Lai 2023).
+An alternative scenario involves secular resonances with the massive circumstellar disk, which can excite planetary eccentricities up to $\sim$0.6, even for modest disk eccentricities (Li and Lai 2023).
 Kozai–Lidov oscillations induced by a distant third body, such as the stellar companion GJ 3305 AB, can also generate large eccentricities (Kozai 1962; Lidov 1962).
@@
 2025), such a companion would likely be Jovian or sub-Jovian ($\lesssim 1$$2,M_{\mathrm{Jup}}$).
-Continued astrometric and radial velocity monitoring will further constrain this scenario (Lacour et al.
+Continued astrometric and RV monitoring will further constrain this scenario (Lacour et al.
 2021).
@@
 To highlight these correlated structures, it is convenient to compute the autocorrelation function (ACF) of the CCF.
-Fig.
-4 presents the individual ACFs of the 4 CCFs.
+Figure 4 presents the individual ACFs of the 4 CCFs.
 The first three epochs are very consistent.
@@
 However, this is beyond the scope of the present paper and we consider the $v\sin i$ estimate of the fourth epoch to be unreliable.
-Figure 3: Power-spectral density (PSD) of the data filtered from the stellar speckles estimate of our four epochs.
-The PSD was lowpass filtered in order to focus on its low-frequency content.
+Figure 3: PSD of the data filtered from the stellar speckles estimate of our four epochs.
+The PSD was lowpass-filtered in order to focus on its low-frequency content.
 A secondary axis is displayed at the top, which corresponds to an equivalent velocity.
@@
 This peak is particularly prominent in the data from the last epoch, when the planet was closer to the host star.
-Figure 4: Normalized Autocorrelation Functions (ACF) of the individual CCFs.
+Figure 4: Normalized ACFs of the individual CCFs.
 The first three epochs show highly consistent ACFs, whereas the last epoch shows deeply correlated structures in the CCF at multiples of 11 km/s.
@@
 Appendix C Method for computing the cross-correlation functions
-To compute the cross-correlation functions, we use the formalism defined in Zucker (2003). The cross-correlation function is defined as:
+To compute the cross-correlation functions, we use the formalism defined in Zucker (2003). The cross-correlation function is defined as
 $CCF(RV)=\frac{\sum_{i}d_{HF}(i)t_{HF}^{RV}(i)}{\sqrt{\sum_{i}d_{HF}(i)^{2}}\sqrt{\sum_{i}t_{HF}^{RV}(i)^{2}}}.$ (1)
 […]
-Figure 5: Individual CCF for the 4 epochs.
+Figure 5: Individual CCFs for the four epochs.
 Appendix D Orbit posterior distribution
-Fig.
-6 presents the posterior distribution of the orbital parameters of 51 Eri b, without our RV measurements (in red) and with our RV measurements (in blue).
+Figure 6 presents the posterior distribution of the orbital parameters of 51 Eri b, without our RV measurements (in red) and with our RV measurements (in blue).
 The inclusion of the HiRISE RV measurements brings significant improvements to the orbital parameters.
@@
 $C=f(\phi)\,A\,\left(\frac{R}{a}\right)^{2},$ (2)
-with $R$ the radius of the planet, $a$ its semi-major axis, $A$ its albedo and $f$ the phase function.
-We considered 3 different values for the phase function to show the impact of the phase of the planet.
+with $R$ the radius of the planet, $a$ its semi-major axis, $A$ its albedo, and $f$ the phase function.
+We considered three different values for the phase function to show the impact of the phase of the planet.
 For $A$, we considered a value of 0.5, similar to that of Jupiter in visible (Karkoschka 1998).
@@
 In the Kozai-Lidov scenario, the orbit of a planet is perturbed by a distant third body, which can lead to oscillations in the inclination and eccentricity of the inner companion.
-The timescale of such oscillations is on the order of magnitude of (Ford and Rasio 2008):
+The timescale of such oscillations is on the order of magnitude of (Ford and Rasio 2008)
 $P_{\mathrm{Kozai}}\approx P_{\mathrm{planet}}\frac{M_{\star}}{M_{\mathrm{pert}}}\left(\frac{a_{\mathrm{pert}}}{a_{\mathrm{planet}}}\right)^{3}\left(1-e_{\mathrm{pert}}^{2}\right)^{3/2}.$ (3)
@@
 We computed the period of the Kozai-Lidov oscillations for different values of $M_{\mathrm{pert}}$, $a_{\mathrm{pert}}$ and $e_{\mathrm{pert}}$ using the posterior distributions of $P_{\mathrm{planet}}$ and $a_{\mathrm{planet}}$ inferred in this study.
-Fig.
-8 shows the results of this analysis.
+Figure 8 shows the results of this analysis.
 The red curve shows the age of the system.
@@
 Additional constrains on the possibility of GJ 3305 AB being in Kozai-Lidov resonance with 51 Eri b could be brought by a direct measurement of the RV of the stellar companion, as it would help constrain both the inclination and the longitude of ascending node of the stellar companion.
-The Kozai-Lidov resonance requires the mutual inclination I between the 2 objects in resonance satisfies (Kozai 1962; Lidov 1962):
+The Kozai-Lidov resonance requires the mutual inclination I between the 2 objects in resonance satisfies (Kozai 1962; Lidov 1962)
 $39\degree<I<141\degree.$ (4)
 From the RV measurement of the stellar companion, we can better constrain its inclination and longitude of ascending node.
-Then, the mutual inclination between 51 Eri b and GJ 3305 AB is computed as (Murray and Dermott 2000):
+Then, the mutual inclination between 51 Eri b and GJ 3305 AB is computed as (Murray and Dermott 2000)
 $\cos I=\cos i_{1}\cos i_{2}+\sin i_{1}\sin i_{2}\cos(\Omega_{1}-\Omega_{2}),$ (5)
-with i1, i2, $\Omega_{1}$, $\Omega_{2}$ the inclinations and longitude of ascending node of the 2 companions.
+with i1, i2, $\Omega_{1}$, $\Omega_{2}$ the inclinations and longitude of ascending node of the two companions.